Half the measure of the angle is 25 more than one third the measure of its supplement
Answer by checkley77(12569)

x/2=25+1/3(180-x/2)
x/2=25+180/3-x/2*1/3
x/2=25+60-x/6
x/2+x/6=85
(6x+2x)/12=85
8x=1,020
x=1,20/8
x=127.5 ans.
180-127.5=52.5 ans.
Proof:
127.5/2=25+1/3(180-127.5/2)
63.75=25+60-127.5/6
63.75=85-21.25
63.75=63.75